Solve the equation:
1000(2^(2/x)) = 3450
Thank you

Answer by Alan3354(69443)

1000(2^(2/x)) = 3450
2^(2/x) = 3.45
(2/x)*log(2) = log(3.45)
2/x = log(3.45)/log(2)
x/2 = log(2)/log(3.45
x = log(4)/log(3.45)